OpenStudy (anonymous):

How to do 4x + 5y = 12 ?

OpenStudy (anonymous):

is this for a linear equation?

OpenStudy (anonymous):

or no

OpenStudy (anonymous):

Yeah it is. I have to solve for Y, but i dont really get it.

OpenStudy (anonymous):

wait are you solving for y or is it y=mx+b

OpenStudy (anonymous):

Oh sorry i meant y=mx+b

OpenStudy (anonymous):

ok great. So what i always do is try to get y on its own. Subtract 4x on the left and add it(in negative form) on the right, so you have 5y=4x+12. Then, keeping isolating y in mind, divide the entire equation by 5, which gives you y=4/5x+12/5. That's really it. If you have more questions like these message me @cooldude8484 and i will be happy to help.

OpenStudy (anonymous):

@gaswimmer4life do you know the swim club meadowbrook in maryland?

OpenStudy (anonymous):

Yeah Ive heard of them, they have a really big pool! @cooldude8484

OpenStudy (anonymous):

just wondering because there was a kinda big meet there a few months ago and i was there with my gf and we saw a georgia team so i was just wondering if you were on it or not

OpenStudy (anonymous):

Oh cool! I am on the Calhoun Blue Barracudas, I dont think my team was there but it would be a lot of fun

OpenStudy (anonymous):

@cooldude8484

OpenStudy (shamil98):

4x + 5y = 12 5y = -4x + 12 y = -4/5x + 12/5

OpenStudy (shamil98):

slope is -4/5 , y intercept is 12/5.

OpenStudy (anonymous):

sorry @gaswimmer4life was trying to find what team it was

OpenStudy (anonymous):

haha wow @shamil98 kinda already answered it...

OpenStudy (anonymous):

Oh haha thats cool its a small team because its a really small town but im pretty sure we have a website @cooldude8484

OpenStudy (anonymous):

@shamil98 Thank you

OpenStudy (anonymous):

ok maybe i'll look u guys up :) @gaswimmer4life

OpenStudy (anonymous):

Haha thanks lol, its alot of fun :) @cooldude8484